find the equation of the tangent plane and the normal line given the surface xy^2z^3=8 at (2,2,1)


1
Expert's answer
2021-10-25T03:39:12-0400

Given, equation of the surface is "f(x,y,z)=xy^2z^3-8" .

And ,we need to find the equation of tangent plane and normal line at (x0,y0,z0)=(2,2,1).

"\frac{\partial f }{\partial x}=y^2z^3,\frac{\partial f }{\partial y}=2xyz^3,\frac{\partial f }{\partial z}=3xy^2z^2\\
\text{At point (2,2,1),}
\frac{\partial f }{\partial x}=4,
\frac{\partial f }{\partial y}=8,
\frac{\partial f }{\partial z}=24\\
\text{Equation of tangent plane,}\\
\frac{\partial f }{\partial x}(x-x_0)+\frac{\partial f }{\partial y}
(y-y_0)+\frac{\partial f }{\partial z}
(z-z_0)=0\\
4(x-2)+8(y-2)+24(z-1)=0\\
4x+8y+24z=48\\
\text{Equation of normal line,}\\
\frac{x-x_0}{\frac{\partial f }{\partial x}}=\frac{y-y_0}{\frac{\partial f }{\partial y}}=\frac{z-z_0}{\frac{\partial f }{\partial z}}\\
\frac{x-2}{4}=\frac{y-2}{8}=\frac{z-1}{24}"